# Optimal map-making with singularities

### 11 Jan 2024 v2

@@ 1 Introduction
The precise measurement of the temperature and polarization of the Cosmic Microwave Background (CMB) has allowed us to establish the standard cosmological model ($\Lambda$CDM). The theory of inflation was developed in the 1980s to address the flatness, horizon, and monopole concerns within the $\Lambda$CDM framework [1, 2]. Despite inflationary cosmology’s enormous success, the expected primordial gravitational waves (PGW) have not been discovered yet. PGW detection and inflationary cosmology testing remain critical parts of current and future CMB researches. The most promising technique to detect the PGW is via the CMB B-mode polarization observations, according to [3, 4, 5].
With the completion of the Planck mission [6], the fundamental scientific goals of various present and future CMB studies are set to detecting the CMB B-mode polarization. These include the ground-based experiments like BICEP/Keck [7, 8], SPTpol [9], CLASS[10], ACTpol [11], POLARBEAR [12], and AliCPT [13, 14, 15] in the Northern Hemisphere; as well as the balloon experiments like EBEX [16], SPIDER [17] and future experiments like SO [18], CMB-S4 [19], and the space mission LiteBIRD [20], which is scheduled to be launched in 2029.
@@ 2.1 The classical minimum variance solution and its limitations
    Each row of $\bm{A}$ contains only one element that is equal to 1, and all other elements are zero. Thus, $\bm{A}^{t}\bm{A}=\bm{N}_{obs}$ is diagonal [^1] Otherwise at least one row will contain more than one elements that are equal to 1 and contains the number of observations at each sky pixel.
@@
    The noise covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ is invertible and fully represents the noise property.
@@
Certainly, in instances where the aforementioned conditions are not satisfied, particularly regarding items 2 and 3, the above solution is no longer applicable. Especially, when the singularity exists in either or both of $\bm{A}^{t}\bm{A}$ and $\bm{C}$, the true optimal solution can be complicated, which has not been documented in detail before.
@@ 2.2 The general minimum variance solution with singularity
First of all, it is possible to obtain a general true optimal solution, which requires to use the singular value decomposition to carefully analyze the linear system and the origin of singularity, in order to clearly separate the singularity that arise from various origins, and use all available information properly to obtain the final optimal solution. Details of the true optimal solution can be found in appendix A.1 – A.3, and the main process includes first dealing with the covariance matrix’s singularity, and then treating the main matrix’s singularity properly, to obtain the final optimal solution by eq. (A.101).
@@ 3.1 The noise amplification effect in optimal mapmaking
On the contrary, although the same problem may also exist in the inversion of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ (the solution with modified pseudo inverse), the chance to get a very small eigenvalue of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ is greatly reduced, because we have eliminated all zero modes of $\bm{C}$ via the modified pseudo inverse, which make the solution with modified pseudo inverse much more robust.
@@ 4 Code validation, simulation and tests
In ground-based CMB experiments, the methods for converting TOD into maps primarily involve the naive map-making, as used by BICEP [30, 31], POLARBEAR [32], and SPT [33]); and also the maximum likelihood map-making, as employed by ACT [34, 35] and POLARBEAR [26]). While naive map-making is prone to producing imperfect estimates of the sky signal, it can be improved by fine-tuning the filters to match the data property, which enables a more diagonal time-domain noise covariance matrix after filtering. Next, signal-only simulations can be used to correct the angular power spectrum suppression cause by filtering [30, 32, 33]. On the other hand, maximum likelihood map-making, particularly as used by POLARBEAR [26], produces maps with a better signal-to-noise ratio by simultaneously solving for the intended signal and contamination templates, and ACT processes the filtered TOD similarly to unfiltered data [34], treating filters as having little effect on the sky signal.
@@ 4.1 The performance tests: for a large sky region
wherein $\nu$ represents the frequency while $\sigma_{0}^{2}$ is the variance of white noise, and $f_{\rm{knee}}$ is the knee frequency. According to the Kolmogorov model of turbulence, the power-law index of the integrated atmospheric emission (noise) varies between $-11/3$ and $-8/3$, thus here we set $\alpha=-3$. For other parameters, we considered an equivalent detector that the noise has been averaged by multi-detectors. Thus, we set the amplitude of white noise $\sigma_{0}^{2}=10^{2}~{}\mu{\rm K}^{2}~{}{\rm Hz}^{-1}$ while the knee frequency $f_{\rm knee}=2.15~{}{\rm Hz}$. When the frequency is less than 0.01 Hz, $P(\nu)$ is constant, which aims to avoid the singularity at $\nu=0$. The power spectrum density for both CMB and atmospheric noise are shown in Figure 1. As elucidated in Section 3.2, the atmospheric emission simulation employed in this study neglects the spatial correlation among the scan rings, leading to a simplification of the noise covariance. It is crucial to emphasize that the noise model we have utilized solely takes into account temporal correlations, while disregarding spatial correlations, as previously discussed in section 3.2.
@@
In the context of a singular covariance matrix $\bm{C}$, the map-level outcomes are depicted in Figure 2, wherein the residual of the $\bm{S}_{\times}$ solution is visually lower compared to the other two solutions.
@@
The angular power spectra of the pixel domain results are derived from an ensemble of 400 simulations and given in Figure 3, which shows that, after eliminating the average noise spectrum and correcting the suppression effect arising from filtering (pertaining exclusively to the $\bm{S}_{+}$ solution) and masking (pertaining to all solutions), unbiased angular power spectra can be obtained from all solutions. Notably, the $\bm{S}_{\times}$ solution displays significantly reduced error bar amplitudes in comparison to both the $\bm{S}_{+}$ solution (for $\ell<80$) and the $\bm{S}_{\rm c}$ solution (for $\ell>30$). Therefore, the $\bm{S}_{\times}$ solution outperforms the other two solutions across the majority of $\ell$-ranges.
Figure 3: The angular power spectrum results corresponding to Figure 2, with the mean noise spectra removed and the suppression factors at each $\ell$ corrected. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
@@ 4.2 The performance tests: for a small sky region
In an actual observational experiments, the scanning strategy is not circular. Instead, the telescope scanned back-and-forth at a fixed elevation angle. In this case, we pay more attention to the spatial correlation of atmospheric emission.
Assuming the atmosphere is a vast 2D plane and the power spectrum of the emission intensity satisfies $P(\bm{k})=\left<f(\bm{k})f^{*}(\bm{k}^{\prime})\right>\propto|\bm{k}|^{-3}\delta(\bm{k}-\bm{k}^{\prime})$, where $f(\bm{k})$ is the 2D Fourier transformation of the atmospheric emission intensity. The covariance in real domain between any two points can be derived trivially:
@@
where $\delta\bm{r}_{ij},\delta t_{ij}$ denotes the physical distance and time difference between two samples, respectively. $\bm{v}_{w}$ denotes the wind speed. For the sake of simplicity, it is treated as a constant in the simulation.
Any linear filter $\bm{M}$ operating in the TOD domain will also distort the covariance matrix. The relationship between the filtered and unfiltered covariance is simply $\tilde{\bm{C}}=\bm{M}\bm{C}\bm{M}^{t}$. Here, we also employ a high-pass filter with the same parameters as in the previous section. The only difference is that the filter operating on each half-scan independently, which means the filter matrix $\bm{M}$ is a block diagonal matrix.
The filtered covariance $\tilde{\bm{C}}$ is singular due to $\bm{M}$ has some zero eigenvalues. The subsequent map-making processing is similar to that mentioned in the previous section. The map-level results are depicted in Figure 4, along with the corresponding residuals for each solution. The angular power spectra results are illustrated in Figure 5.
@@
Figure 5: The angular power spectrum results corresponding to Figure 4, with the mean noise spectra removed and the suppression factors at each $\ell$ corrected. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
@@ 5 Summary and discussions
In this paper, we show how to solve for the sky map in an optimal way with a full consideration of the singularity problems. In summary, the singular value decomposition, along with a detailed investigation of the effective noise covariance matrix, should be employed to discover the general true optimal solution. On the other hand, a sub-optimal solution that is just slightly different from the classical pseudo inverse method, was designed to greatly simplify the true optimal solution. Simulations with the AliCPT observation strategies confirm that the sub-optimal solution is capable of producing significantly better sky maps and angular power spectra than the naive and classical pseudo inverse solutions, especially at the large scales.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    When $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are both singular: Can be one of above three.
However, we need to point out the following fact: as mentioned in section 3.1, when the mapmaking algorithm is designed to keep the signal lossless (as a precondition), there is going to be a significant amplification of the noise term in the modes with poor SNR, and the overall mapmaking result may looks bad. This is an inevitable consequence of all “lossless” or “make lossless at all costs” mapmaking algorithms. The solution is also simple: one needs to consider a posterior Wiener filtering style approach to put a reasonable balance between the signal loss and noise control.
In principle, it is even possible to further augment the method in this work, e.g., to solve for the best estimation of the CMB angular power spectrum directly from the TOD, which means to re-write the main equation into
@@
whose partial derivative to $C_{\ell}$ is easy to obtain, and hence enables the QML calculations. Ideally, this can give the best estimation of $C_{\ell}$ even without map-making; however, because we need to work in the pixel domain for the foreground removal and delensing, such a direct solution of $C_{\ell}$ is not a good choice unless those pixel domain operations can be successfully transported to the time domain, which is an exciting idea but does not seems to be feasible in the near future.
The concept of the near-optimal solution can readily assist in the solution of singularity issues in other methods, such as the maximum likelihood solution, so long as the method calls for the inversion of a matrix with a singularity. Simply use the modified pseudo-inverse for the inversion of the noise covariance matrix and the standard pseudo-inverse for all other matrix inversions. The benefits of the near-optimal solution presented in this study are then fully realized.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China No. 2021YFC2203100 and No. 2021YFC2203104, the Anhui Provincial Natural Science Foundation 2308085MA30 and the Anhui project Z010118169.
@@ Appendix A The mathematical details
A.1 How to deal with the noise covariance matrix’s singularity
@@
Because the covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ does converge, it can always be decomposed into $\bm{C}=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$, where $\bm{P}$ is unitary and $\bm{\lambda}$ is diagonal (allow zero diagonal elements). Therefore, it is always possible to convert $\bm{n}$ into a standard white noise $\bm{n}_{0}$ that follows the standard normal distribution and satisfy $\langle\bm{n}_{0}\bm{n}_{0}^{t}\rangle=\bm{I}$. The connection is given by $\bm{n}=\bm{P}\bm{\lambda}\bm{n}_{0}$, which automatically gives $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$; thus, we get
@@
Because $\bm{\lambda}$ may have zero diagonal element, we divide $\bm{\lambda}$ and $\bm{n}_{0}$ into blocks as
@@
Apparently, this equation should first be solved for the lower part because it effectively contains no noise:
@@
Let $\bm{x}_{1}=\bm{G}^{t}\bm{x}$ be a column vector, whose first $\ell$ elements are denoted as $\bm{x}_{1,upper}$, and the rest, $\bm{x}_{1,lower}$. We write $\bm{P}^{t}_{2}\bm{A}\bm{x}$ explicitly as:
@@
which is a precise solution (noiseless). Now we can return to eq. (A.20) to deal with its upper half: $\bm{d}_{1}=\bm{P}_{1}^{t}\bm{A}\bm{x}+\bm{\lambda}_{1}\bm{n}_{1}$. Substitute $\bm{x}=\bm{G}\bm{x}_{1}$, we get
@@
$\bm{d}_{1}-\bm{A}_{1,left}\bm{x}_{1,\rm{upper}}=\bm{d}_{1}^{\prime}=\bm{A}_{1,\rm{right}}\bm{x}_{1,\rm{lower}}+\bm{\lambda}_{1}\bm{n}_{1}.$ (A.41)
@@
A.2 How to safely reduce the linear system’s size

### 28 Feb 2024 v3

@@ 1 Introduction
The precise measurement of the temperature and polarization of the Cosmic Microwave Background (CMB) has allowed us to establish the standard cosmological model ($\Lambda$CDM). The theory of inflation was developed in the 1980s to address the flatness, horizon, and monopole concerns within the $\Lambda$CDM framework [1, 2]. Despite inflationary cosmology’s enormous success, the expected primordial gravitational waves (PGW) have not been discovered yet. Detection of PGW and testing inflationary cosmology remain critical parts of current and future CMB researches. The most promising technique to detect the PGW is via the CMB B-mode polarization observations, according to [3, 4, 5].
With the completion of the Planck mission [6], the fundamental scientific goals of various present and future CMB studies are set towards detecting the CMB B-mode polarization. These include the ground-based experiments like BICEP/Keck [7, 8], SPTpol [9], CLASS[10], ACTpol [11], POLARBEAR [12], and AliCPT [13, 14, 15] in the Northern Hemisphere; as well as the balloon experiments like EBEX [16], SPIDER [17] and future experiments like SO [18], CMB-S4 [19], and the space mission LiteBIRD [20], which is scheduled to be launched in 2029.
@@ 2.1 The classical minimum variance solution and its limitations
    Each row of $\bm{A}$ contains only one element that is equal to 1, and all other elements are zero. Thus, $\bm{A}^{t}\bm{A}=\bm{N}_{obs}$ is diagonal [^1] Otherwise at least one row will contain more than one element equal to 1 and contains the number of observations at each sky pixel.
@@
    The noise covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ is invertible and fully represents the noise properties.
@@
Certainly, in instances where the aforementioned conditions are not satisfied, particularly regarding items 2 and 3, the above solution is no longer applicable. Especially, when a singularity exists in either or both of $\bm{A}^{t}\bm{A}$ and $\bm{C}$, the true optimal solution can be complicated, which has not been documented in detail before.
@@ 2.2 The general minimum variance solution with singularity
First of all, it is possible to obtain a general true optimal solution, which requires using the singular value decomposition to carefully analyze the linear system and the origin of singularity, in order to clearly separate the singularity that arise from various origins, and use all available information properly to obtain the final optimal solution. Details of the true optimal solution can be found in appendix A.1 – A.3, and the main process includes first dealing with the covariance matrix’s singularity, and then treating the main matrix’s singularity properly, to obtain the final optimal solution by eq. (A.101).
@@ 3.1 The noise amplification effect in optimal mapmaking
On the contrary, although the same problem may also exist in the inversion of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ (the solution with modified pseudo inverse), the chance to get a very small eigenvalue of $\bm{A}^{t}\bm{C}^{\times}\bm{A}$ is greatly reduced, because we have eliminated all zero modes of $\bm{C}$ via the modified pseudo inverse, which makes the solution with modified pseudo inverse much more robust.
@@ 4 Code validation, simulation and tests
In ground-based CMB experiments, the methods for converting TOD into maps primarily involve naive map-making, as used by BICEP [30, 31], POLARBEAR [32], and SPT [33]); and also maximum likelihood map-making, as employed by ACT [34, 35] and POLARBEAR [26]). While naive map-making is prone to producing imperfect estimates of the sky signal, it can be improved by fine-tuning the filters to match the data property, which enables a more diagonal time-domain noise covariance matrix after filtering. Next, signal-only simulations can be used to correct the angular power spectrum suppression cause by filtering [30, 32, 33]. On the other hand, maximum likelihood map-making, particularly as used by POLARBEAR [26], produces maps with a better signal-to-noise ratio by simultaneously solving for the intended signal and contamination templates, and ACT processes the filtered TOD similarly to unfiltered data [34], treating filters as having little effect on the sky signal.
@@ 4.1 The performance tests: for a large sky region
wherein $\nu$ represents the frequency while $\sigma_{0}^{2}$ is the variance of white noise, and $f_{\rm{knee}}$ is the knee frequency. According to the Kolmogorov model of turbulence, the power-law index of the integrated atmospheric emission (noise) varies between $-11/3$ and $-8/3$, thus here we set $\alpha=-3$. For other parameters, we considered an equivalent detector of which the noise level is same to the noise level that is averaged over all detectors. Thus, we set the amplitude of white noise $\sigma_{0}^{2}=10^{2}~{}\mu{\rm K}^{2}~{}{\rm Hz}^{-1}$ while the knee frequency $f_{\rm knee}=2.15~{}{\rm Hz}$. When the frequency is less than 0.01 Hz, $P(\nu)$ is constant, which aims to avoid the singularity at $\nu=0$. The power spectrum density for both CMB and atmospheric noise are shown in Figure 1. As elucidated in Section 3.2, the atmospheric emission simulation employed in this study neglects the spatial correlation among the scan rings, leading to a simplification of the noise covariance. It is crucial to emphasize that the noise model we have utilized solely takes into account temporal correlations, while disregarding spatial correlations, as previously discussed in section 3.2.
@@
In the context of a singular covariance matrix $\bm{C}$, the map-level outcomes are depicted in Figure 2. The figure shows the maps reconstructed by the solutions $\bm{S}_{\rm c},\bm{S}_{\times}$ and $\bm{S}_{+}$, as well as their respective residuals from left to right. It is visually apparent that the recommended solution $\bm{S}_{\times}$ (lower-middle) has lower residual compared to the other two solutions (lower-left and lower-right).
@@
Considering the map-making procedure as a form of linear transformation for TOD, we generated $N=400$ sets of CMB-only and noise-only simulations, with the observed sky map for the $i$-th set given by
$\bm{x}_{i}^{\rm obs}=\bm{x}_{i}^{\rm CMB}+\bm{x}_{i}^{\rm noise}.$ (4.5)
And the noise-debiased power spectrum for the $i$-th observation maps $\bm{x}^{\rm obs}_{i}$ is
$\hat{C}_{\ell,i}=\hat{C}^{\rm obs}_{\ell,i}-\frac{1}{N-1}\sum_{k\neq i}\hat{C}^{\rm noise}_{\ell,k}$ (4.6)
The angular power spectrum $\hat{C}_{\ell}$ is estimated from the pseudo power spectrum $\tilde{C}_{\ell}$ by the MASTER method [38] with the NaMaster implementation [39], where the full sky angular power spectrum is recovered from the partial sky spectrum with an $\ell\times\ell$ coupling matrix, determined by the sky mask with optional apodization. In this work, the apodization is done with a 3-degree C2 configuration, and the basic angular power spectrum computation is done using the anafast routine in the HEALPix package [40].
The average and variance of the angular power spectra with bin width $\Delta\ell=20$ of the solutions $\bm{S}_{c},\bm{S}_{\times}$ and $\bm{S}_{+}$ are depicted in Figure 3 using green, red, and blue solid lines, respectively; which shows that, after eliminating the average noise spectrum and correcting the suppression effect arising from filtering (pertaining exclusively to the $\bm{S}_{+}$ solution) and masking (pertaining to all solutions), unbiased angular power spectra can be obtained from all solutions. Notably, the $\bm{S}_{\times}$ solution displays significantly reduced error bar amplitudes in comparison to both the $\bm{S}_{+}$ solution (for $\ell<80$) and the $\bm{S}_{\rm c}$ solution (for $\ell>30$). Therefore, the $\bm{S}_{\times}$ solution outperforms the other two solutions across the majority of $\ell$-ranges.
Figure 3: The angular power spectrum results corresponding to Figure 2, with the mean noise spectra removed and the suppression factors at each bin corrected. The bin width is $\Delta\ell=20$ for $61010\leq\ell<$ 610. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
@@ 4.2 The performance tests: for a small sky region
In actual observational experiments, the scanning strategy is not circular. Instead, the telescope scans back-and-forth at a fixed elevation angle. In this case, we pay more attention to the spatial correlation of atmospheric emission.
Assuming the atmosphere is a vast 2D plane and the power spectrum of the emission intensity satisfies $P(\bm{k})=\left<f(\bm{k})f^{*}(\bm{k}^{\prime})\right>\propto|\bm{k}|^{-3}\delta(\bm{k}-\bm{k}^{\prime})$, where $f(\bm{k})$ is the 2D Fourier transformation of the atmospheric emission intensity, the covariance in real domain between any two points can be derived trivially:
@@
where $\delta\bm{r}_{ij},\delta t_{ij}$ denote the physical distance and time difference between two samples, respectively. $\bm{v}_{w}$ denotes the wind speed. For the sake of simplicity, it is treated as a constant in the simulation.
Any linear filter $\bm{M}$ operating in the TOD domain will also distort the covariance matrix. The relationship between the filtered and unfiltered covariance is simply $\tilde{\bm{C}}=\bm{M}\bm{C}\bm{M}^{t}$. Here, we also employ a high-pass filter with the same parameters as in the previous section. The only difference is that the filter operates on each half-scan independently, which means the filter matrix $\bm{M}$ is a block diagonal matrix.
The subsequent map-making processing for the case of a small sky region is similar to that mentioned in the previous section. Figure 4 displays the solutions $\bm{S}_{c},\bm{S}_{\times}$, and $\bm{S}_{+}$ with their residuals from left to right at the map-level. Similar to Figure 2 and even more evident, the $\bm{S}_{\times}$ solution has the lowest residual (lower-middle), while the residual of $\bm{S}_{c}$ (lower-left) is dominated by instrumental noise and the residual of $\bm{S}_{+}$ (lower-right) is dominated by large-scale CMB loss due to filtering. The angular power spectra results corresponding to Figure 4 are illustrated in Figure 5. It is notable that the solution $\bm{S}_{\times}$ achieves the best error bar across almost all angular scales.
@@
Figure 5: The angular power spectrum results corresponding to Figure 4, with the mean noise spectra removed and the suppression factors at each $\ell$ corrected. The bin width is $\Delta\ell=20$ for $31010\leq\ell<$ 310. Top panel: the corrected spectra of one realization as example. Middle panel: difference between the expected CMB angular power spectrum and each solutions’ mean spectrum of 400 simulations. Bottom panel: the corresponding error bar amplitudes at each $\bm{\ell}$.
@@ 5 Summary and discussions
In this paper, we show how to solve for the sky map in an optimal way with a full consideration of the singularity problems. In summary, to get the true optimal solution, the singular value decomposition should be employed along with a detailed investigation of the effective noise covariance matrix, which is the most complex part of the solution. Once this is done as indicated in the appendix, the rest of the solution can be done with a regular singular-value decomposition to discover the general true optimal solution. This approach also tells us, that the general true optimal solution depends on the situation of singularity, and cannot be achieve with just one simple equation.
Considering the complexity of the true optimal solution, we have also designed a simplified sub-optimal solution that can be given in just one simple equation, which is only slightly different from the classical pseudo inverse method and is hence much easier to implement. The side effect of this sub-optimal solution is strictly limited to the singular part, which is usually negligible. Moreover, this sub-optimal solution also helps to greatly suppress the unwanted noise amplification effect, as discussed in section 3.1. Simulations with the AliCPT observation strategies confirm that this sub-optimal solution is capable of producing significantly better sky maps as well as angular power spectra than the naive and classical pseudo inverse solutions, especially at the large scales.
As a possible augment of the method in this work, one may even solve for the best estimation of the CMB angular power spectrum directly from the TOD, which means to re-write the main equation into
@@
whose partial derivative to $C_{\ell}$ is easy to obtain, and hence enables the QML calculations. Ideally, this can give the best estimation of $C_{\ell}$ even without map-making; however, because we need to work in the pixel domain for the foreground removal and delensing, such a direct solution of $C_{\ell}$ is not a good choice unless those pixel domain operations can be successfully transported to the time domain, which is an exciting idea but does not seems to be feasible in the near future. However, we still aim to test a low-resolution version of this idea as an interesting follow-up of this work.
The concept of the near-optimal solution can readily assist in the solution of singularity issues in other approaches, such as the maximum likelihood solution, so long as the method calls for the inversion of a singular covariance matrix. Despite the difficulty of obtaining the true optimal solution of the singularity problem, it is easy to incorporate the recommended sub-optimal solution in other approaches: when one needs to invert a singular noise covariance matrix, the inversion should done with a modified pseudo-inverse as explained in appendix A.5.2, and other matrix inversions can be done with the regular inverse or pseudo-inverse (if necessary). By this simple operation, the benefits of the recommended solution is fully realized, even in another approach.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China No. 2021YFC2203100 and No. 2021YFC2203104, the Anhui Provincial Natural Science Foundation 2308085MA30 and the Anhui project Z010118169. We thank the anonymous referee for the useful comments and James Creswell for helping improve this work.
@@ Appendix A The mathematical details
A.1 How to deal with the singularity of the noise covariance matrix
@@
Because the covariance matrix $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle$ does converge, it can always be decomposed into $\bm{C}=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$, where $\bm{P}$ is unitary and $\bm{\lambda}$ is diagonal (allowing zero diagonal elements). Therefore, it is always possible to convert $\bm{n}$ into a standard white noise $\bm{n}_{0}$ that follows the standard normal distribution and satisfy $\langle\bm{n}_{0}\bm{n}_{0}^{t}\rangle=\bm{I}$. The connection is given by $\bm{n}=\bm{P}\bm{\lambda}\bm{n}_{0}$, which automatically gives $\bm{C}=\langle\bm{n}\bm{n}^{t}\rangle=\bm{P}\bm{\lambda}^{2}\bm{P}^{t}$; thus, we get
@@
Because $\bm{\lambda}$ may have zero diagonal elements, we divide $\bm{\lambda}$ and $\bm{n}_{0}$ into blocks as
@@
Evidently, this equation should first be solved for the lower part because it effectively contains no noise:
@@
Let $\bm{x}_{1}=\bm{G}^{t}\bm{x}$ be a column vector, whose first $\ell$ elements are denoted as $\bm{x}_{1,\rm{upper}}$, and the rest, $\bm{x}_{1,\rm{lower}}$. We write $\bm{P}^{t}_{2}\bm{A}\bm{x}$ explicitly as:
@@
which is a precise solution (noiseless). Now we can return to eq. (A.20) to deal with its upper half: $\bm{d}_{1}=\bm{P}_{1}^{t}\bm{A}\bm{x}+\bm{\lambda}_{1}\bm{n}_{1}$. Substituting $\bm{x}=\bm{G}\bm{x}_{1}$, we get
@@
$\bm{d}_{1}-\bm{A}_{1,\rm{left}}\bm{x}_{1,\rm{upper}}=\bm{d}_{1}^{\prime}=\bm{A}_{1,\rm{right}}\bm{x}_{1,\rm{lower}}+\bm{\lambda}_{1}\bm{n}_{1}.$ (A.41)
@@
A.2 How to safely reduce the size of the linear system